# How does the entanglement entropy of a many-body quantum system change after a single measurement?

### v2

@@ Abstract
For one-dimensional free Dirac fermions, we compute numerically the probability distribution of the change in the entanglement entropy (EE), after the saturation time, resulting from a single measurement of the occupation number by using different measurement protocols. For the quantum jump and the projective measurement protocols, we observe clear deviations from Gaussianity characterized by broader and asymmetric tails, exponential for positive values of the change, and a peak at zero that increases with the system size and the monitoring strength supporting that in all cases the EE is in the area law phase. Another distinct feature of the distribution is its spatial inhomogeneity. In the weak monitoring limit, the distribution is close to Gaussian with a broad support for boundary point separating the two subsystems defining the EE while for the rest of sites has asymmetric exponential tails and a much narrower support. For a quantum state diffusion protocol, the distribution is Gaussian for weak monitoring. As the monitoring strength increases, it gradually develops symmetric exponential tails. In the strong monitoring limit, the tails are still exponential but the core turns from Gaussian to strongly peaked at zero suggesting the dominance of quantum Zeno effect. For all monitoring strengths, the distribution is size independent.
@@ I Introduction
The description of the impact of measurements on the quantum dynamics has been a recurrent research topic from the early days [1] of quantum theory. Indeed, one of the accepted postulates of quantum mechanics [2] is that, after a measurement, the wavefunction collapses to one eigenstate of the operator represented by the observable being measured with a probability given by the Born’s rule, namely, the module squared of the overlap between the state before the measurement and the mentioned eigenstate. However, these so-called projective measurements are only a small subset of the experimental protocols [3, 4] that are currently available to gain information on a quantum system. Experimental [5] and theoretical developments [6, 7, 8, 6, 9] in quantum optics in the last three decades has broadened enormously the horizon of quantum measurements. Typical examples beyond projective measurement are continuous measurements based on photo-detection, where the continuous monitoring is interrupted by quantum jumps. [10, 11, 12, 5, 13]. This quantum jump protocol led to the development of the quantum trajectories formalism [7, 8, 6, 9] to model theoretically the experimental results which is numerically more efficient, and has a sharper physical interpretation, than the use of the full Lindbladian formalism [14, 15] typical of open quantum systems. Another popular continuous measurement protocol is the one based on homodyne detection [16, 17, 4, 18], where at each time step the system is perturbed by a weak measurement characterized by a Gaussian noise [19, 20, 21, 22].
@@
The quantum measurement problem has gained new impetus in recent times not only because the availability of the mentioned experimental measurement set-ups with a high level of control but also because for quantum computation, and other quantum technologies, it is fundamental to extract information from quantum states with a minimal level of disturbance on the system so that quantum coherence is preserved. This is a challenging task as repeated measurements can induce [23, 24, 25, 26, 27, 21, 28, 29, 30, 31, 32, 33, 34] measurement induced phase transitions (MIPT) in the entanglement entropy, and other quantum information observables non-linear in the density matrix, separating phases with different scaling with respect to the system size. Therefore, a high monitoring rate has the potential of destroying quantum features that may be necessary for practical applications. Although challenging because of the post-selection problem, the existence of MIPT has already [35, 36] been confirmed experimentally.
The precise characterization of the MIPT in many-body quantum systems is still poorly understood. The main features are still being debated even for one dimensional free fermions with no disorder. By mapping the using the replica trick and mapping the problem to a non-linear sigma model [37, 38, 39, 40], it has been reported [41] that no transition occurs for Dirac fermions subjected to projective measurements. However, a transition has been predicted in Refs. [42, 20] based on a combination of numerical results for relatively small sizes and semi-analytic arguments. For Majorana fermions, and a continuous monitoring protocol based on weak measurements, the use of a variant of the mentioned non-linear sigma model predicts [43, 44] the existence of a transition at finite monitoring rate. Increasing the spatial dimensionality [45], the range of hopping [46, 47] or including interactions [48, 49, 50, 51] seems to induce a MIPT at finite monitoring strength. The interplay between Anderson localization induced by disorder and measurements promises a rich phase diagram which only recently has started to be investigated [52, 53].
@@
In the context of the Sachdev-Ye-Kitaev (SYK) model [54, 55, 56, 57, 58] a one dimensional chain with an integrable SYK on each site was shown [59] to experience a volume to area transition analytically for sufficiently large monitoring rate by using the Keldysh path-integral formalism. In the gravity dual, monitoring, at least phenomenologically and in the post-select approximation, seems to be related with additional complex sources in the boundary [60, 61]. By contrast, a two-site SYK initially uncoupled but being continuously monitored undergoes [62] a transition to a state whose gravity dual is a traversable wormhole [63] provided that the monitoring rate is sufficiently strong.
@@ II Model and Method
where $c_{i}$ and $c_{i}^{\dagger}$, $i=1,2,\cdots L$ are the annihilation and creation operators respectively for fermions at site $i$. $J$ is the hopping strength which for convenience is set to $J=1$. We impose periodic boundary conditions $c_{i}=c_{i+L},c_{i}^{\dagger}=c_{i+L}^{\dagger}$ with $L$ the lattice size which we set to be even. The Hamiltonian commutes with the particle number operator $\hat{N}=\sum_{i=1}^{L}\hat{n}_{i}$ where $\hat{n}_{i}\equiv c^{\dagger}_{i}c_{i}$ is the occupation number operator at site $i$ which has eigenvalues $0,1$. The eigenvalues of $\hat{N}$ are the total number of fermions $N=1,2,\ldots$. We consider the system at half-filling. The initial state, termed Néel state, in an occupation number basis is $|\psi(t=0)\rangle=|101010\cdots\rangle$.
@@
We consider the measurement of the occupation number at site $i$, $\hat{n}_{i}$ by three different protocols:
@@
(1). Quantum state diffusion (QSD) protocol [19, 20, 21, 22], also termed homodyne detection [16, 17, 4, 18] or weak measurements, is characterized by coupling the quantum system to a stochastic environment. The time evolution of the wave-function is governed by an Itô-type stochastic Schrödinger equation with a weak Gaussian noise, parametrized by $\gamma$. Further details of this protocol are provided in Appendix  A.1.
@@
(2). Quantum jump (QJ) protocol [10, 11, 12, 5, 13], a continuous monitoring protocol usually referred to as photo-detection in quantum optics where the measurement apparatus is always active and the detection, usually termed a quantum jump, occurs with a certain probability that depends on the monitoring strength $\gamma$. In our case, what is detected is an occupied site, $n_{i}=1$. We model the time evolution of the state by the quantum trajectory method [7, 8, 6, 9] characterized by the jump operator $L_{i}=\sqrt{\gamma}\hat{n}_{i}$. Further details are in Appendix  A.2.
@@
(3). Projective measurement (PM) protocol refers to the standard description of measurements in quantum mechanics [2] where the observable to be measured is represented by a Hermitian operator. As a consequence of the measurement, the wavefunction of the system collapses to one eigenstate of this operator with a probability given by the Born’s rule, namely, the absolute value of the overlap between the eigenstate and the wavefunction before the measurement. In our case, we measure the occupation number at site $i$ represented by the operator $\hat{n}_{i}$ which could have two outcome $0,1$. The measuring rate per site, controlled by a Poisson distribution, is given by $\gamma$. Further details are found in Appendix  A.3.
@@
A crucial feature that simplifies enormously the numerical treatment of the model is that since $\hat{n}_{i}^{2}=\hat{n}_{i}$, the operator controlling the state evolution, even if subjected to measurements of the occupation number, will still be quadratic and preserve the particle number. As a consequence, generic states $|\psi(t)\rangle$ are Gaussian so they can be written down as [19],
@@
Strictly speaking, rather than with the distribution of the full EE at those long times, our main interest is the distribution of the EE difference between the state right before and right after a measurement. However, for the sake of completeness, we depict in Fig. 2 the probability distribution of the full EE in this long time region. As is observed, in the weak monitoring limit, the distribution is close to Gaussian for the three protocols. Deviations from Gaussianity, qualitatively similar in all protocols, require a strong monitoring rate corresponding to a region where the quantum Zeno effect [74, 75, 76, 77, 78] is dominant. Therefore, the distribution function of the full EE seems to be sensitive to the existence of a transition in the EE due to measurements. We stress that our system size $L=512$ is too small to rule out that for weak monitoring rate, the distribution of the EE in the $L\to\infty$ limit will still be Gaussian. However, it is promising that within the range of sizes we can explore, the full distribution captures sharp differences between the weak and strong monitoring rates.
We now turn to the study of the distribution before and after a measurement which is the main interest of this paper. We initiate our analysis with the QSD protocol.
@@ III The quantum state diffusion measurement protocol
We recall that the details of the QSD protocol are given in Appendix A.1. Since we are mostly interested in the effect of measurement in on the EE, we define the speed of the entanglement change at each step $dt$ of the dynamical evolution as
@@
Figure 3: The probability distribution of $\delta S_{qsd}$ for the QSD protocol Eq. (8), employing Eq. (7) after the saturation time of the average EE. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 3 to 3. The distribution agrees well with a Gaussian distribution in the weak monitoring limit. For intermediate monitoring strength, the distribution is well described by a simple function interpolating from Gaussian, for small $\delta S_{qsd}$ to exponential, for large $\delta S_{qsd}$. In the strong monitoring limit, the exponential tails are still observed. However, for small $\delta S_{qsd}$ , the distribution is strongly peaked around zero which indicates the dominance of quantum Zeno effect and provides an upper bound for the area law in the EE. We do not observe any size dependence in our results, see the text for an explanation of this feature.
Finally, we address an important feature of the distribution $P(\delta S_{qsd})$: its independence on the system size $L$. We observe this feature for all monitoring strengths $\gamma$ and in the full range of sizes we can explore numerically so it is likely a robust feature of the protocol not related to limitations in the maximum size we can reach numerically. Heuristically, this can be understood by a scaling analysis that exploits the fact that, for this protocol only, all sites are equally and simultaneously monitored at each time step. As a consequence of this, and the stochastic nature of the dynamics, $\delta S_{qsd}$ is also the summation over $\ell$ Gaussian random numbers of average zero and variance $O(1/L)$, so the resulting distribution will be a Gaussian of average zero and size independent variance provided that the subsystem size $\ell$ scales with the total size $L$. Therefore the distribution $P(\delta S_{qsd})$ should also be Gaussian which is consistent with our numerical results.
We note that the observed size independence and Gaussianity of $\delta S_{qsd}$ is not related to the size dependence of the entanglement entropy itself. Therefore, at least for small $\gamma$, we cannot make any statement about the existence of the area law by looking at the distribution of $\delta S_{qsd}$. For that purpose, it would be necessary to map the model close to the Fermi energy onto a non-linear sigma model and follow the procedure of Ref. [41] for the PM protocol. Given that qualitative similarities between both protocols, including the use of Dirac fermions, we believe that it is likely that the EE will also follow an area law for any finite monitoring rate.
IV The QJ measurement protocol
@@
where $\tau^{-}$ is the time right before the quantum jump, $S_{A}$ is computed using Eq. (7), and $\tau^{+}$ the time right after the quantum jump.
@@
The rest of features of the distribution are also different from those of the QSD protocol. The tails of $P(\Delta S_{qj})$ are asymmetric for all monitoring strength. For $\Delta S_{qj}>0$, the tail is exponential for any $\gamma$ if we neglect the effect of the peak at zero and, for strong monitoring, the region $\Delta S_{qj}\geq 0.5$ which decays faster but has a comparatively small weight. The existence of states so that $\Delta S_{qj}>0$ indicates that, counterintuitively, in some cases measurements characterized by quantum jumps can increase the EE, this feature was earlier observed in Ref. [64]. By contrast, the left tail is more sensitive to the monitoring strength. We could not find a simple analytic expression for general $\gamma$ though for weak monitoring, putting aside the effect of the peak at zero, is well described by a Gaussian though the agreement seems to become worse as the system size increases.
Figure 4: The probability distribution of the change in entanglement entropy $\Delta S_{qj}$ Eq. (9), computed employing Eq. (7), due to a quantum jump in the QJ protocol, for sufficiently late times so that EE has reached the saturation value. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 4 to 4.
@@
Figure 5: $\Delta S_{qj}$ Eq. (9) as a function of the measured sites $r=1,\ldots L$. The left plot is for the weak monitoring rate $\gamma=0.1$ and system size $L=1024$, while the right plot is for the strong monitoring rate $\gamma=3$ and the system size is $L=512$. Note the stark differences between boundary points, separating the two subsystems defining the EE, and the rest.
An explicit calculation of the distribution function separating sites close to the boundary from those in the bulk, see Fig. 6, confirms the qualitative picture above. For weak monitoring, the distribution function for the boundary point is a Gaussian with a small peak at zero. As more sites around the boundary are added, the distribution develops a narrow exponential tail for $\Delta S_{qj}\gtrsim 0$. We note that rare instances of $\Delta S_{qj}\gtrsim 0$ occur due to the fact that the measurement not only disentangles the target site but also rearranges the remaining particles across the rest of the $L-1$ sites, effectively enhancing entanglement in unmeasured regions. This counterintuitive effect highlights the non-local feedback inherent in quantum measurement dynamics. For bulk sites, the distribution is asymmetric with exponential tails characterized by different decay rates. For strong monitoring, the distribution is effectively a delta function at zero for bulk points (note the scale of the axis), while for boundary points, though there is also a peak at zero, we observe an asymmetric distribution with a broad support that covers all possible values of $\Delta S_{qj}$.
@@
Figure 6: The distribution function of $\Delta S_{qj}$ Eq. (9) for specific sites. The upper panel: the monitoring rates is $\gamma=0.1$ and the system size is $L=1024$. The bottom panel: the monitoring rates is $\gamma=3.0$ and the system size is $L=512$. Here sites 1, 2 and 3 stand for sites around the boundary of the subsystem. Site 4 stands for the site at the center of the subsystem. Specifically, sites 1 are sites $[r=1,r=L/2,r=L/2+1,r=L]$, sites 2 are $[r=1+5,r=L/2-5,r=L/2+6,r=L-5]$, sites 3 are $[r=1+15,r=L/2-15,r=L/2+16,r=L-15]$ and sites 4 $[r=L/4,r=3L/4]$. Note that for Fig. 6, the scale of $\Delta S_{qj}$ is of the order $10^{-10}$ and the maximum is $10^{10}$ so it is effectively a Dirac delta function.
Figure 7: The density-map of $\Delta S_{qj}$ as a function of the occupation number of the selected site that are going to be measured. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 7 to 7. The system size for the upper panel is $L=1024$, and the lower panel is $L=512$.
In order to gain a more comprehensive understanding of the effect of the quantum jumps on the EE, we study its dependence on the occupation number $n_{i}$ at site $i$ right before the measurement occurs. For that purpose, we show in Fig. 7, the density-map of $\Delta S_{qj}$ as a function of $n_{i}$. We note that in the QJ protocol only the outcome $n_{i}=1$ is measured. It is expected [64] that the smaller the occupation number, the larger the change in the EE after a measurement. From the results in Fig. 5, sufficiently large changes are restricted to the boundary of the subsystem. As a result, a change in $n_{i}=0\rightarrow 1$ near the subsystem boundary, induced by a measurement, is expected to generate a more pronounced change to the EE. In the weak monitoring limit, most sites will have a filling rate near $1/2$ with virtually no sites with a filling rate near 00, so that we cannot observe the expected large entanglement change for sites with $n_{i}\sim 0$. Instead, we observe the expected behavior of a maximum at half filling and, consistent with previous results for the distribution of $\Delta S_{qj}$, a relatively strong peak at $\Delta S_{qj}=0$. It is also worth to noting that even in this weak monitoring limit, there is already a small peak at $n_{i}=1$ as a consequence of the measurements. Indeed, even for intermediate monitoring strength, the peak rapidly moves to $n_{i}=1$ which illustrate the strong impact of measurement on the entanglement dynamics.
In the strong monitoring limit, $n_{i}$ can take any value though the change in EE has a cut-off value $\ln{2}\approx 0.69$, reflecting the maximal entanglement contribution from a single particle. As mentioned earlier, this maximum value can only occur at the subsystem boundary sites because for the rest of sites measurements induce little change in $\Delta S_{qj}$. Therefore, a measurement would only affect particles around the boundary.

### v3

@@ Abstract
For one-dimensional non-interacting complex fermions, we compute numerically the probability distribution of the change in the entanglement entropy (EE), after saturation, resulting from a single measurement of the occupation number by using different measurements protocols. In the thermodynamic limit, the system is in the area-law phase for any monitoring strength, however we can also study the distribution in the critical phase characterized by a logarithmic scaling of the EE with system size by considering a sufficiently weak monitoring strength. For the quantum jump and the projective measurement protocols, we observe clear deviations from Gaussianity characterized by broader and asymmetric tails, exponential for positive values of the change, and a peak at zero, likely a precursor of Zeno effect, that increases with the system size and the monitoring strength. Intriguingly, the distribution is spatially inhomogeneous. For sites around the boundary separating the subsystems defining the EE, the distribution is close to Gaussian with a broad support while for the rest of sites has asymmetric exponential tails and a much narrower support. As the monitoring strength increases, the full distribution is controlled by the boundary sites. For a quantum state diffusion protocol, where the change in EE is defined between two consecutive time steps, the distribution, Gaussian for weak monitoring, gradually develops symmetric exponential tails. For strong monitoring, the core turns from Gaussian to strongly peaked at zero suggesting the dominance of quantum Zeno effect.
@@ I Introduction
The description of the impact of measurements on the quantum dynamics has been a recurrent research topic from the early days [1] of quantum theory. Indeed, one of the accepted postulates of quantum mechanics [2] is that, after a measurement, the wavefunction collapses to one eigenstate of the operator represented by the observable being measured with a probability given by the Born’s rule, namely, the module squared of the overlap between the state before the measurement and the mentioned eigenstate. However, these so-called projective measurements are only a small subset of the experimental protocols [3, 4] that are currently available to gain information on a quantum system. Experimental [5] and theoretical developments [6, 7, 8, 6, 9] in quantum optics in the last three decades has broadened enormously the horizon of quantum measurements. Typical examples beyond projective measurement are continuous measurements based on photodetection, where the continuous monitoring is interrupted by quantum jumps. [10, 11, 12, 5, 13]. This quantum jump protocol led to the development of the quantum trajectories formalism [7, 8, 6, 9] to model theoretically the experimental results which is numerically more efficient, and has a sharper physical interpretation, than the use of the full Lindbladian formalism [14, 15] typical of open quantum systems. Another popular continuous measurement protocol is the one based on homodyne detection [16, 17, 4, 18], where at each time step the system is perturbed by a weak measurement characterized by a Gaussian noise [19, 20, 21, 22].
@@
The quantum measurement problem has gained new impetus in recent times not only because the availability of the mentioned experimental measurement set-ups with a high level of control but also because for quantum computation, and other quantum technologies, it is fundamental to extract information from quantum states with a minimal level of disturbance on the system so that quantum coherence is preserved. This is a challenging task as repeated measurements can trigger [23, 24, 25, 26, 27, 21, 28, 29, 30, 31, 32, 33, 34] a measurement induced phase transitions (MIPT) in the entanglement entropy, and other quantum information observables non-linear in the density matrix, separating phases with different scaling with respect to the system size. Therefore, a high monitoring rate has the potential of destroying quantum features that may be necessary for practical applications. Although challenging because of the post-selection problem, the existence of MIPT has already [35, 36] been confirmed experimentally.
The precise characterization of the MIPT in interacting many-body quantum systems are still being debated. However, for one dimensional free Dirac fermions, namely, non-relativistic complex fermions, with no disorder and subjected to either continuous [19] or projective [37] measurements there is now growing consensus, despite some initial contradicting results [38, 20], that the system stays in the area-law phase for any monitoring strength. For projective measurements [37], see also [39, 40], this conclusion was reached analytically by using the replica trick and mapping the problem onto a non-linear sigma model [41, 42, 43, 44].
For Majorana fermions, and a continuous monitoring protocol based on weak measurements, the use of a variant of the mentioned non-linear sigma model predicts [45, 40] the existence of a transition at finite monitoring rate. The different symmetries of the associated non-linear sigma model explains the differences between Majorana and Dirac fermions.
Increasing the spatial dimensionality [46], the range of hopping [47, 48, 49] or including interactions [50, 51, 49, 52, 53, 54, 55] seems to induce a MIPT at finite monitoring strength. The interplay between Anderson localization induced by disorder and measurements promises a rich phase diagram which only recently has started to be investigated [56, 57].
@@
In the context of the Sachdev-Ye-Kitaev (SYK) model [58, 59, 60, 61, 62] a setting consisting in two one dimensional coupled chains with an integrable SYK on each site, where the parity is monitored, was shown [63] to experience a MIPT transition analytically for sufficiently large monitoring rate by using the Keldysh path-integral formalism. In the gravity dual, monitoring, at least phenomenologically and in the post-select approximation, seems to be related with additional complex sources in the boundary [64, 65]. By contrast, a two-site SYK initially uncoupled but being continuously monitored undergoes [66] a transition to a state whose gravity dual is a traversable wormhole [67] provided that the monitoring rate is sufficiently strong.
@@ II Model and Method
where $c_{i}$ and $c_{i}^{\dagger}$, $i=1,2,\cdots L$ are the annihilation and creation operators respectively for fermions at site $i$. $J$ is the hopping strength which for convenience is set to $J=1$. We impose periodic boundary conditions $c_{i}=c_{i+L},c_{i}^{\dagger}=c_{i+L}^{\dagger}$ with $L$ the lattice size which we set to be even. The Hamiltonian commutes with the particle number operator $\hat{N}=\sum_{i=1}^{L}\hat{n}_{i}$ where $\hat{n}_{i}\equiv c^{\dagger}_{i}c_{i}$ is the occupation number operator at site $i$ which has eigenvalues $0,1$. The eigenvalues of $\hat{N}$ are the total number of fermions $N=1,2,\ldots$. We consider the system at half-filling $N=L/2$. The initial state, termed Néel state, in an occupation number basis is $|\psi(t=0)\rangle=|101010\cdots\rangle$.
@@
We consider the measurement of the occupation number at site $i$, $\hat{n}_{i}$, by three different protocols:
@@
(1). Quantum state diffusion (QSD) protocol [19, 20, 21, 22], also termed homodyne detection [16, 17, 4, 18] or weak measurements, is characterized by coupling the quantum system to a stochastic environment. The time evolution of the wave-function is governed by an Itô-type stochastic Schrödinger equation with a weak Gaussian noise, parametrized by $\gamma$. Further details of this protocol are provided in Appendix A.1.
@@
(2). Quantum jump (QJ) protocol [10, 11, 12, 5, 13], a continuous monitoring protocol usually referred to as photo-detection in quantum optics where the measurement apparatus is always active and the detection, usually termed a quantum jump, occurs with a certain probability that depends on the monitoring strength $\gamma$. In our case, what is detected is an occupied site, $n_{i}=1$. We model the time evolution of the state by the quantum trajectory method [7, 8, 6, 9] characterized by the jump operator $L_{i}=\sqrt{\gamma}\hat{n}_{i}$. Further details are in Appendix A.2.
@@
(3). Projective measurement (PM) protocol refers to the standard description of measurements in quantum mechanics [2] where the observable to be measured is represented by a Hermitian operator. As a consequence of the measurement, the wavefunction of the system collapses to one eigenstate of this operator with a probability given by the Born’s rule, namely, the absolute value of the overlap between the eigenstate and the wavefunction before the measurement. In our case, we measure the occupation number at site $i$ represented by the operator $\hat{n}_{i}$ which could have two outcomes $0,1$. The measuring rate per site, controlled by a Poisson distribution, is given by $\gamma$. Further details are found in Appendix A.3.
@@
A crucial feature that simplifies enormously the numerical treatment of the model for all these protocols is that since $\hat{n}_{i}^{2}=\hat{n}_{i}$, the operator controlling the state evolution, even if subjected to measurements of the occupation number, will still be quadratic and preserve the particle number. As a consequence, generic states $|\psi(t)\rangle$ are Gaussian so they can be written down as [19],
@@
Strictly speaking, rather than with the distribution of the full EE at those long times, our main interest is the distribution of the EE difference between the state right before and right after a measurement – for homodyne detection this corresponds to one time step. However, for the sake of completeness, we depict in Fig. 2 the probability distribution of the full EE in this long time region. As is observed, in the weak monitoring limit, the distribution is close to Gaussian for the three protocols. Deviations from Gaussianity, characterized by broad tails and a bimodal distribution are qualitatively similar in all protocols, require a strong monitoring rate which we believe corresponds to a region where the quantum Zeno effect [78, 79, 80, 81, 82] is dominant. We comment that similar results for the full EE distribution were obtained earlier [31] for a continuously monitored Ising chain. These results suggest that the distribution function of the full EE seems to be capable of capturing a transition in the EE (if present). As we mentioned in the introduction, this model is in the area-law phase for $L\to\infty$, and any $\gamma>0$, but we still believe that in the range of $L$ that we explore, and for small $\gamma$, our results describe the distribution of a system in a phase where the EE scales with the logarithmic of the system size. In any case, it is promising that within the range of sizes we can explore, the full distribution captures sharp differences between the weak and strong monitoring rates.
We now turn to the study of the distribution before and after a measurement which is the main interest of this paper.
@@ III The quantum state diffusion measurement protocol
We initiate our analysis with continuous measurements characterized by the QSD protocol, see Appendix A.1 for details. This protocol, related to homodyne measurements, is characterized by weak measurements of the particle number on each site of the lattice at each time step $dt$. Therefore, by change in entanglement after one measurement, we really mean the change in EE after one time step because, by definition, at a given time, measurements are carried out on all sites at once so the change after a single measurement is not possible. It is natural to study the speed of the entanglement change at each step $dt$ of the dynamical evolution defined as,
@@
Figure 3: The probability distribution of $\delta S_{qsd}$ for the QSD protocol Eq. (8), employing Eq. (7) after the saturation time of the average EE. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 3 to 3. The distribution agrees well with a Gaussian distribution in the weak monitoring limit. For intermediate monitoring strength, the distribution is well described by a simple function interpolating from Gaussian, for small $\delta S_{qsd}$, to exponential, for large $\delta S_{qsd}$. In the strong monitoring limit, the exponential tails are still observed. However, for small $\delta S_{qsd}$ , the distribution is strongly peaked around zero which indicates the dominance of quantum Zeno effect and provides an upper bound for the area law in the EE. We do not observe any size dependence in our results, see the text for an explanation of this feature.
Finally, we address an important feature of the distribution $P(\delta S_{qsd})$: its independence on the system size $L$. We observe this feature for all monitoring strengths $\gamma$ and in the full range of sizes we can explore numerically so it is likely a robust feature of the protocol not related to limitations in the maximum size we can reach numerically. Although we do not have a clear explanation of this scale invariance, we believe that the reason is that the change in the EE is governed by the measurement of sites around the boundary between the two subsystems which does not scale with $L$. This is a feature in the QJ and PM protocols, see Fig. 5, which we think it is at work for the QSD protocol as well. Since all sites are measured simultaneously, it is difficult to differentiate the role of each site without changing the protocol.
We note that the observed size independence and Gaussianity of $\delta S_{qsd}$ is not related to the size dependence of the entanglement entropy itself. Therefore, at least for small $\gamma$, we cannot make any statement about the existence of the area law for $L\to\infty$ by looking at the distribution of $\delta S_{\rm qsd}$. For that purpose, it would be necessary to map the model onto a non-linear sigma model, following the procedure of Ref. [37], for the PM protocol. Indeed, it has been recently shown [39, 83, 84] that other measurement protocols the EE still verifies an area law at any finite monitoring rate provided that the symmetry of the model does not change.
IV The quantum jump measurement protocol
@@
where $S_{A}$ is computed using Eq. (7) and $\tau^{\pm}$ the time right before (-) and right after (++) the quantum jump.
@@
The rest of features of the distribution are also different from those of the QSD protocol. The tails of $P(\Delta S_{qj})$ collapse after rescaling them by the system size $L$, but they are still asymmetric for all monitoring strength. For $\Delta S_{qj}>0$, the tail is exponential for any $\gamma$ if we neglect the effect of the peak at zero and, for strong monitoring, the region $\Delta S_{qj}\geq 0.5$ which decays faster but has a comparatively small weight. The existence of states so that $\Delta S_{qj}>0$ indicates that, counterintuitively, in some cases measurements characterized by quantum jumps can increase the EE, this feature was earlier observed in Ref. [68]. By contrast, the left tail is more sensitive to the monitoring strength. We could not find a simple analytic expression for general $\gamma$, yet for weak monitoring, putting aside the effect of the peak at zero, the distribution is well described by a Gaussian though the agreement seems to become worse as the system size increases.
Figure 4: The probability distribution of the change in entanglement entropy $\Delta S_{qj}$ Eq. (9), computed employing Eq. (7), due to a quantum jump in the QJ protocol, for sufficiently late times so that EE has reached the saturation value. The $P(\Delta S_{qj})$ is rescaled by multiplying the system size $L$. The tail is independent of system size after rescaling, signaling the tail is dominated by the boundary region. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 4 to 4.
@@
Figure 5: $\Delta S_{qj}$ Eq. (9) as a function of the measured sites $r=1,\ldots,L$. The left plot is for the weak monitoring rate $\gamma=0.1$ and system size $L=1024$, while the right plot is for the strong monitoring rate $\gamma=3$ and the system size is $L=512$. Note the stark differences between boundary points, separating the two subsystems defining the EE, and the rest.
Therefore, we expect that when the subsystem size is much larger than this boundary region, $P(\Delta S_{qj})$ would be independent of the subsystem size $\ell$, which is illustrated in Fig. 6. In the weak monitoring limit $\gamma=0.1$, when the subsystem size is smaller than $128$, $P(\Delta S_{qj})$ changes substantially from a broader distribution to a narrower distribution and finally become insensitive to $\ell$. When the monitoring is stronger, $\gamma\geq 0.5$, the boundary region is much smaller than subsystem size $\ell\geq 16$ that we investigate, so $P(\Delta S_{qj})$ do not depend on the subsystem size. For the projective measurement protocol, we also obtain a similar results (not shown).
Figure 6: The probability distribution $P(\Delta S_{qj})$ for different subsystem sizes $\ell$. Only in the weak measurement limit $\gamma=0.1$, $P(\Delta S_{qj})$ depends on $\ell$ due to finite size effects. For stronger monitoring, $P(\Delta S_{qj})$ is insensitive to $\ell$ which may be indicative of an area-law behavior. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 6 to 6. The system size is $L=512$.
An explicit calculation of the distribution function separating sites close to the boundary from those in the bulk, see Fig. 7, confirms the qualitative picture above. For weak monitoring, the distribution function for the boundary point is a Gaussian with a small peak at zero. As more sites around the boundary are added, the distribution develops a narrow exponential tail for $\Delta S_{qj}\gtrsim 0$. We note that rare instances of $\Delta S_{qj}\gtrsim 0$ occur due to the fact that the measurement not only disentangles the target site but also rearranges the remaining particles across the rest of the L1L-1 sites, effectively enhancing entanglement in unmeasured regions. This counterintuitive effect highlights the non-local feedback inherent in quantum measurement dynamics. For bulk sites, the distribution is asymmetric with exponential tails characterized by different decay rates. For strong monitoring, the distribution is effectively a delta function at zero for bulk points (note the scale of the axis), while for boundary points, though there is also a peak at zero, we observe an asymmetric distribution with a broad support that covers all possible values of $\Delta S_{qj}$.
@@
Figure 7: The distribution function of $\Delta S_{qj}$ Eq. (9) for specific sites. The upper panel: the monitoring rates is $\gamma=0.1$ and the system size is $L=1024$. The bottom panel: the monitoring rates is $\gamma=3.0$ and the system size is $L=512$. Here sites 1, 2 and 3 stand for sites around the boundary of the subsystem. Site 4 stands for the site at the center of the subsystem. Specifically, sites 1 are sites $[r=1,r=L/2,r=L/2+1,r=L]$, sites 2 are [r=1+5,r=L/25,r=L/2+6,r=L5][r=1+5,r=L/2-5,r=L/2+6,r=L-5], sites 3 are [r=1+15,r=L/215,r=L/2+16,r=L15][r=1+15,r=L/2-15,r=L/2+16,r=L-15] and sites 4 $[r=L/4,r=3L/4]$. Note that for Fig. 7, the scale of $\Delta S_{qj}$ is of the order $10^{-10}$ and the maximum is $10^{10}$ so it is effectively a Dirac delta function.
Figure 8: The density-map of $\Delta S_{qj}$ as a function of the occupation number of the selected site that are going to be measured. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 8 to 8. The system size for the upper panel is $L=1024$, and the lower panel is $L=512$.
In order to gain a more comprehensive understanding of the effect of the quantum jumps on the EE, we study its dependence on the occupation number $n_{i}$ at site $i$ right before the measurement occurs. We show in Fig. 8, the density-map of $\Delta S_{qj}$ as a function of $n_{i}$. We note that in the QJ protocol only the outcome $n_{i}=1$ is measured. It is expected [68] that the smaller the occupation number, the larger the change in the EE after a measurement. From the results in Fig. 5, sufficiently large changes are restricted to the boundary of the subsystem. As a result, a change in $n_{i}=0\rightarrow 1$ near the subsystem boundary, induced by a measurement, is expected to generate a more pronounced change to the EE. In the weak monitoring limit, most sites will have a filling rate near $1/2$ with virtually no sites with a filling rate near 0, so that we cannot observe the expected large entanglement change for sites with $n_{i}\sim 0$. Instead, we observe the expected behavior of a maximum at half filling and, consistent with previous results for the distribution of $\Delta S_{qj}$, that a relatively strong peak occurs at $\Delta S_{qj}=0$. It is also worth noting that even in this weak monitoring limit, there is already a small peak at $n_{i}=1$ as a consequence of the measurements. Indeed, even for intermediate monitoring strength, the peak rapidly moves to $n_{i}=1$ which illustrate the strong impact of measurement on the entanglement dynamics.
In the strong monitoring limit, $n_{i}$ can take any value though the change in EE has a cut-off value $\ln{2}\approx 0.69$, reflecting the maximal entanglement contribution from a single particle. As mentioned earlier, this maximum value can only occur at the subsystem boundary sites because, for the rest of sites, measurements induce little change in $\Delta S_{qj}$. Therefore, a measurement would only affect particles around the boundary.